# Simple fits for the neutrino luminosities from protoneutron star cooling

### v2

@@ I Introduction
The neutrino signal from a future Galactic supernova (SN) explosion represents one of the next frontiers of neutrino astrophysics (see, e.g., Ref. [1]). Existing and planned large underground neutrino detectors guarantee that a high-statistics neutrino burst will be collected during such an event (see, e.g., Ref. [2] for a review). This detection will be extremely important to probe the SN explosion mechanism [3], neutrino flavor conversions [4], and particle physics [5] occurring in the deepest stellar regions.
In order to characterize the response of a detector to a SN neutrino burst one has to rely on the outcome of state-of-art numerical SN simulations. Different from the case of solar neutrinos, a standard prediction for SN neutrino fluxes does not exist. Therefore, many current studies employ a parametric approach with a suitable range of variation guided by the results of the simulations. For this purpose it has been shown that the time-integrated SN neutrino fluxes are well represented by a spectrum with the functional form [6, 7]
@@
Concerning the time evolution of the neutrino luminosities, most of the detector characterizations use, for comparison, results obtained from numerical tables provided by simulations (see, e.g., [8, 9, 10]). However, in order to extend parametric studies also to the temporal evolution of the neutrino signal, it would be useful to have a simple functional prescription also for the luminosities, based on a few parameters that permit to cover the range of variation obtained in the numerical SN simulations.
In this respect, the seminal work of Loredo and Lamb [11], concerned with an analysis of SN 1987A data, assumed an exponential cooling model. This simple recipe was widely used (see, e.g., Ref. [12]). However, as noticed already in Ref. [13], protoneutron star (PNS) cooling calculations do not yield exponential neutrino light curves but instead suggest that the neutrino luminosity is better described by a power-law decline. A further step was taken in the recent work of Ref. [14], where on grounds of an analysis of recent spherically symmetric PNS cooling simulations covering post-bounce times up to $t\sim 70$ s, a combined ansatz of lower-law and exponential behavior of the form $L_{\nu_{i}}(t)\propto t^{-1}\,e^{-(t/\tau)^{n}}$ was proposed for the long-time behavior of the neutrino light curve, where $\tau$ is a characteristic time scale of the PNS cooling. In that work $\tau\sim 30$ s was adopted to handle the transition to the regime of neutrino transparency, and therefore the functional description of the neutrino light curve follows a simple decline according to $t^{-1}$ in the first 10 s after core bounce. However, the models considered in Ref. [14] did not include the effect of convection in the PNS. This has important consequences for the shape of the neutrino light curve, because the Kelvin-Helmholtz neutrino cooling of the PNS is strongly accelerated if convection persists for many seconds.
@@
Inspired by these previous findings, we perform a detailed analysis of the neutrino luminosities obtained in a set of spherically symmetric PNS cooling models recently described in Ref. [16]. These simulations were performed with six-species neutrino-transport based on a fully energy-dependent two-moment scheme with a variable Eddington closure obtained from the solution of the Boltzmann transport equation. Moreover, the effects of PNS convection and muons were taken into account. After this analysis we propose an accurate fit for the neutrino luminosities during the PNS cooling evolution at post-bounce times $t\gtrsim 1$ s, based on the following functional form:
@@
In the present paper we report our analysis and fitting procedure and provide extended tables with the parameter values for the fit functions of all our models and for interesting correlations between these fit parameters. The outline of the paper is as follows. In Section II we show the neutrino luminosities from our 1D PNS cooling simulations for different PNS masses and nuclear EoSs. In Section III we describe our analytical fitting of the neutrino and antineutrino luminosities of all flavors based on Eq. (2) for the models discussed before. In Section IV we discuss the dependence of the values of the fitting parameters on PNS mass and nuclear EoS and how their determination can help in deducing information on these latter quantities from a neutrino measurement. In Section V we discuss tight correlations of the parameter values on the one hand with the PNS mass (for fixed EoS) and on the other hand between each other (for fixed PNS mass). Then, in Section VI, we consider how the previous results change for models that neglect the effects of PNS convection and of muons in the PNS cooling models. In Section VII we show how our analytical recipe can be used to fit the event rate in a large neutrino underground detector. Finally, in Section VIII we discuss possible applications of our fitting formula and finish with conclusions. In several appendices we provide extended tables with the values obtained by our numerical fitting and needed for practical use.
@@
Table 1: Final times $t_{\rm fin}$ and ratios $X_{\nu_{i}}^{\rm fin}$ [see Eq. (3)] for all PNS simulations and neutrino species. Values of $X_{\nu_{i}}^{\rm fin}$ larger than 0.15 are printed in bold and asterisks mark the largest values, $X_{\nu_{\mu}}^{\rm fin}=0.219$ and $X_{\nu_{\mu}}^{\rm fin}=0.185$, for models 1.62-SFHx-c and for 1.61-LS220-c, respectively.
Refer to captionRefer to caption
@@
II Neutrino luminosities
Our work is based on results from 1D hydrodynamical PNS cooling simulations employing the PROMETHEUS-VERTEX neutrino-hydrodynamics code [18], which solves the fully energy and velocity dependent neutrino transport for all six species of neutrinos and antineutrinos with a state-of-the-art implementation of the neutrino interactions [19, 20, 21]. The models we will use are available in Ref. [22] and are the same as those considered in the recent study of Ref. [16], where we refer the interested readers to for more information on the model evolution and discussion of the neutrino emission properties.
@@ II Neutrino luminosities
We consider simulations that yield baryonic PNS masses [^1] Since the neutrino emission is tightly correlated with properties of the PNS, we use this quantity as reference for a systematic variation instead of the progenitor mass, which is not monotonically related with the PNS mass. of 1.36 $M_{\odot}$, 1.44 $M_{\odot}$, 1.62 $M_{\odot}$, 1.77 $M_{\odot}$, and 1.93 $M_{\odot}$, in each case computed with four different nuclear EoSs, namely DD2 [23, 24, 25], SFHo, SFHx [26, 24], and LS220 with a nuclear incompressibility at saturation density of $K=220$ MeV [27]. Correspondingly, we follow Ref. [16] in our naming convention of the simulations, specifying the PNS mass and the EoS in the model names, e.g. 1.62-DD2. Our standard simulations include PNS convection and muons; those without convection are denoted by a suffix “-c” to their names, and those without muons by a suffix “-m”. In Tab. 1 we list all of the discussed simulations and the final post-bounce times $t_{\rm fin}$ when they were stopped.
@@
Figure 1 displays the time evolution of the luminosities $L_{\nu}(t)$ for all the neutrino (left panels) and antineutrino (right panels) flavors in units of $\mathrm{B/s}$ ($1~{}\operatorname{B}=10^{51}~{}\operatorname{erg}$) in the time interval [1,10] s for the 1.62 $M_{\odot}$ models, computed with the set of considered EoSs cases: DD2 in blue, SFHo in orange, SFHx in green, and LS220 in red. From this figure it is evident that the neutrino luminosities exhibit a change of their slope in the time interval considered. Since we expect a power-law behavior of the neutrino luminosities in the early cooling phase and we also want to enhance the visibility of the change in the steepness of the luminosity decline at later times, we take inspiration from Ref. [14] and show in Fig. 2 the product of the post-bounce time and the luminosities, $t\,L_{\nu}(t)$, for all of the neutrino (left panels) and antineutrino (right panels) flavors in the same time interval and for the same simulations as in Fig. 1. One witnesses an interesting difference of the results for the employed EoSs. For DD2, SFHo and SFHx EoSs and all of the neutrino species the quantity $t\,L_{\nu}$ increases as a power-law at $t\gtrsim 1~{}\operatorname{s}$ to peak around $t\approx 4~{}\operatorname{s}$, followed by a steep suppression at later times. This behavior produces a prominent “knee” at about 5–6 s after bounce. In contrast, for LS220 $t\,L_{\nu}$ reaches its peak earlier, namely at $t\lesssim 3~{}\operatorname{s}$, and it features a slower decline later on. Since for each model the luminosities of all heavy-lepton neutrino species show a similar evolution, we will focus exclusively on the $\bar{\nu}_{\mu}$ signal as representative of all non-electron neutrino species, unless otherwise noted.
Refer to captionRefer to caption
@@
For each case shown in Figs. 2, 3, and 4, the quantity $t\,L_{\nu}$ has a peak at $t_{\nu_{i},\,\rm max}\lesssim 6$ s, when we define ${L_{\nu_{i},\,\rm max}\equiv L_{\nu_{i}}(t_{\nu_{i}\,\rm max})}$. The corresponding values of $t_{\nu_{i},\,\rm max}$ are listed in Tab. 2 for neutrinos and in Tab. 3 for antineutrinos in Appendix A. In Tab. 1 we provide the values of the quantity
@@
which is a measure for how long our simulations followed the late-time luminosity decline until they were stopped at $t_{\rm fin}$. For our benchmark PNS models (all models including muons and convection) $X_{\nu_{i}}^{\rm fin}\lesssim 0.15$ holds for all neutrinos and antineutrinos (the largest value is $X_{\nu_{e}}^{\rm fin}=0.123$ for 1.62-LS220). In order to adopt a well defined, common final time for all of the neutrino species and simulations when fitting the luminosities in the following, we cut the data at an instant $t_{\nu_{i},\rm c}$ when $X_{\nu_{i}}^{\rm c}\equiv t_{\nu_{i},\rm c}L_{\nu_{i},\rm c}/(t_{\nu_{i},\rm max}L_{\nu_{i},\rm max})=0.15$. In addition to the values of $t_{\nu_{i},\rm max}$ we also provide those of $t_{\nu_{i},{\rm c}}$ for all neutrino species and simulations in Tables 2 and 3 in Appendix A. It can be seen that for a fixed EoS and neutrino species, the values of both of these quantities increase with the PNS mass. In contrast, given the PNS mass and EoS, $t_{\nu_{i},\,\rm max}$ and $t_{\nu_{i},{\rm c}}$ do not vary much among the different kinds of neutrinos and antineutrinos (e.g., for 1.62-DD2 $t_{\nu_{i},{\rm max}}\approx 4\,\operatorname{s}$ and $t_{\nu_{i},{\rm c}}\approx 8\,\operatorname{s}$ for all species). However, at fixed PNS mass, $t_{\nu_{i},\,\rm max}$ and $t_{\nu_{i},{\rm c}}$ feature a strong dependence on the considered EoS. In particular, $t_{\nu_{i},{\rm c}}\lesssim 10\,\operatorname{s}$ for all the benchmark models with DD2, whereas SFHo and SFHx lead to slightly larger values of $t_{\nu_{i},{\rm c}}$, with $t_{\nu_{i},{\rm c}}\lesssim 12.5\,\operatorname{s}$ for all models, and in the cases with LS220 we find the highest values of $t_{\nu_{i},{\rm c}}$, with $t_{\nu_{i},{\rm c}}\lesssim 15\,\operatorname{s}$ for the largest PNS mass.
Refer to captionRefer to captionRefer to captionRefer to caption
@@ III Fit functions
where $C$, $\alpha$, $\tau$ and $n$ are free parameters and $t$ and $\tau$ are measured in seconds. [^2] We define the general parameters as $C$, $\alpha$, $\tau$ and $n$ and we use $C_{\nu_{i}}$, $\alpha_{\nu_{i}}$, $\tau_{\nu_{i}}$ and $n_{\nu_{i}}$ when referring specifically to a neutrino species $\nu_{i}$. In Eq. (2) the parameter $C$ is a normalization constant, $\alpha$ describes the power-law behavior at early cooling time deviating from the simple $t^{-1}$ behavior, $\tau$ is a characteristic cooling time for the exponential drop after the peak in $t\,L_{\nu}$, with $n$ representing the strength of the suppression at late times. As an example, we show for the $1.62~{}M_{\odot}$ model and all EoSs the time evolution of $t\,L_{\nu_{i}}$ from simulations (blue lines) and its best fit (orange) between $t=1$ s and $t_{\nu_{i},{\rm c}}$, for $\nu_{e}$ in Fig. 5, for $\bar{\nu}_{e}$ in Fig. 6 and for $\bar{\nu}_{\mu}$ in Fig. 7. For DD2, SFHo and SFHx the fit well reproduces the results from the simulations in the considered time interval, with best-fit parameters in a similar range. For each EoS, the highest values of the fit parameters $C$ and $\tau$ are obtained for $\bar{\nu}_{\mu}$, while $\nu_{e}$ show the lowest ones. A similar behavior is found for the other PNS masses.
@@
The fit for models with EoS LS220 shows worse agreement with the data from simulations (see the lower-right panels in Figs. 5, 6, and 7). In particular, for all neutrino species the fit overestimates the luminosity at $t\approx 1\,$s, exhibits a lower value of the peak, whereas it well reproduces the luminosity at later times. In this case, the values of the best-fit parameters are in a completely different range compared to those of the models with the other EoSs. In particular, for each neutrino species $C\approx\mathcal{O}(100)~{}\operatorname{B}/\operatorname{s}$ and $\alpha<0$, whereas $\tau$ and $n$ adopt much smaller values than for the other considered EoSs, with $\tau\lesssim 0.1\,$s and $n\lesssim 0.5$, due to the more shallow decline of the luminosity at late times. The differences in the time dependence of the neutrino luminosities between models with different EoSs can be understood by the different behavior of the nuclear symmetry energy as a function of baryon density for the considered EoS. In the LS220 EoS, the symmetry energy exhibits a steeper increase with baryon density than in all other cases. As discussed in Ref. [15], a larger positive derivative of the symmetry energy with baryon density can lead to suppressed convection in the PNS mantle at high densities and low electron fraction. This effect happens in the simulations with the LS220 EoS after about 3 s and gradually quenches PNS mantle convection, thus delaying subsequent PNS neutrino cooling due to reduced neutrino luminosities. In contrast, PNS convection continues to be active in a spatially more extended region including the PNS mantle in simulations with the other three EoSs, for which reason convectively enhanced neutrino luminosities are maintained for a longer period of time until a late, steep decline follows when the PNS has deleptonized and cools off. These differences explain the early kink in the neutrino luminosities for models with the LS220 EoS, whereas a prominent knee-like shape of $L_{\nu}(t)$ can be witnessed for the DD2, SFHo, and SFHx models (see Fig. 1). These differences motivate us to define DD2, SFHo, and SFHx as members of an EoS-class that we call “Class A”, whereas LS220 is a representative of a “Class B.”

### v3

@@ I Introduction
The neutrino signal from a future Galactic supernova (SN) explosion represents one of the next frontiers of neutrino astrophysics (see, e.g., Refs. [1, 2, 3, 4, 5]). Existing and planned large underground neutrino detectors guarantee that a high-statistics neutrino burst will be collected during such an event (see, e.g., Refs. [1, 2, 6, 7, 8, 9, 10, 11]). This detection will be extremely important to probe the SN explosion mechanism [12, 13, 14, 15, 16, 17, 18, 19], neutrino flavor conversions [20, 21, 22, 23, 24], and particle physics [25, 26, 24] occurring in the deepest stellar regions.
In order to characterize the response of a detector to a SN neutrino burst, one has to rely on the outcome of state-of-the-art numerical SN simulations. Different from the case of solar neutrinos, a standard prediction for SN neutrino fluxes does not exist. Therefore, many current studies employ a parametric approach with a suitable range of variation guided by the results of the simulations. For this purpose, it has been shown that the time-integrated SN neutrino fluxes are well represented by a spectrum with the functional form [27, 28]
@@
Concerning the time evolution of the neutrino luminosities, most of the detector characterizations use, for comparison, results obtained from numerical tables provided by simulations (see, e.g., [29, 30, 1, 8, 31, 32]). However, in order to extend parametric studies also to the temporal evolution of the neutrino signal, it would be useful to have a simple functional prescription also for the luminosities, based on a few parameters that permit one to cover the range of variation obtained in the numerical SN simulations.
In this respect, the seminal work of Loredo and Lamb [33], concerned with an analysis of SN 1987A data, assumed an exponential cooling model. [^1] A more sophisticated model accounting for both accretion and cooling phase was presented in Ref. [34]. This simple recipe was often used (see, e.g., Refs. [35, 36, 37]) for schematic estimations. However, as noticed already in Ref. [38], protoneutron star (PNS) cooling calculations do not yield exponential neutrino light curves but instead suggest that the neutrino luminosity is better described by a power-law decline [39]. A further step was taken in the recent work of Ref. [40], where on the grounds of an analysis of recent spherically symmetric PNS cooling simulations covering post-bounce times up to $t\sim 70$ s, a combined ansatz of power-law and exponential behavior of the form $L_{\nu_{i}}(t)\propto t^{-1}\,e^{-(t/\tau)^{n}}$ was proposed for the long-time behavior of the neutrino light curve, where $\tau$ is a characteristic timescale of the PNS cooling.222 Analytical parametrizations of the late-time SN neutrino signal have been also presented in Ref. [41]. In that work $\tau\sim 30$ s was adopted to handle the transition to the regime of neutrino transparency, and therefore the functional description of the neutrino light curve follows a simple decline according to $t^{-1}$ in the first 10 s after core bounce. However, the models considered in Ref. [40] did not include the effect of convection in the PNS. This has important consequences for the shape of the neutrino light curve, because the Kelvin-Helmholtz neutrino cooling of the PNS is strongly accelerated if convection persists for many seconds.
@@
Inspired by these previous findings, we perform a detailed analysis of the neutrino luminosities obtained in a set of spherically symmetric PNS cooling models recently described in Ref. [43]. These simulations were performed with six-species neutrino transport based on a fully energy-dependent two-moment scheme with a variable Eddington closure obtained from the solution of the Boltzmann transport equation. Moreover, the effects of PNS convection and muons were taken into account. After this analysis we propose an accurate fit for the neutrino luminosities during the PNS cooling evolution at post-bounce times $t\gtrsim 1$ s, based on the following functional form:
@@
In the present paper, we report our analysis and fitting procedure and provide extended tables with the parameter values for the fit functions of all our models and for interesting correlations between these fit parameters. The outline of the paper is as follows. In Section II we show the neutrino luminosities from our 1D PNS cooling simulations for different PNS masses and nuclear EoSs. In Section III we describe our analytical fitting of the neutrino and antineutrino luminosities of all flavors based on Eq. (2) for the models discussed before. In Section IV we discuss the dependence of the values of the fitting parameters on PNS mass and nuclear EoS and how their determination can help in deducing information on these latter quantities from a neutrino measurement. In Section V we discuss tight correlations of the parameter values on the one hand with the PNS mass (for fixed EoS) and, on the other hand, between each other (for fixed PNS mass). Then, in Section VI, we consider how the previous results change for models that neglect the effects of PNS convection and of muons in the PNS cooling models. In Section VII we show how our analytical recipe can be used to fit the event rate in a large neutrino underground detector. Finally, in Section VIII we discuss possible applications of our fitting formula and finish with conclusions. In several appendices we provide extended tables with the values obtained by our numerical fitting and needed for practical use.
II Neutrino luminosities
Our work is based on results from 1D hydrodynamical PNS cooling simulations employing the PROMETHEUS-VERTEX neutrino-hydrodynamics code [45], which solves the fully energy and velocity-dependent neutrino transport for all six species of neutrinos and antineutrinos with a state-of-the-art implementation of the neutrino interactions [46, 47, 48]. The models we will use are taken from the Garching Core-collapse Supernova Archive [49] (all data are available upon request) and they are the same as those considered in the recent study of Ref. [43], to which we refer the interested readers for more information on the model evolution and a discussion of the neutrino emission properties.
@@ II Neutrino luminosities
Table 1: Final times $t_{\rm fin}$ and ratios $X_{\nu_{i}}^{\rm fin}$ [see Eq. (3)] for all PNS simulations and neutrino species. Values of $X_{\nu_{i}}^{\rm fin}$ larger than 0.15 are printed in bold and asterisks mark the largest values, $X_{\nu_{\mu}}^{\rm fin}=0.219$ and $X_{\nu_{\mu}}^{\rm fin}=0.185$, for models 1.62-SFHx-c and for 1.61-LS220-c, respectively. The detailed numerical neutrino outputs for all the simulations listed here are available at the Garching Core-Collapse Supernova Archive [49].
@@
We consider simulations that yield baryonic PNS masses [^3] Since the neutrino emission is tightly correlated with properties of the PNS, we use this quantity as reference for a systematic variation instead of the progenitor mass, which is not monotonically related with the PNS mass. of 1.36 $M_{\odot}$, 1.44 $M_{\odot}$, 1.62 $M_{\odot}$, 1.77 $M_{\odot}$, and 1.93 $M_{\odot}$, in each case computed with four different nuclear EoSs, namely DD2 [50, 51, 52], SFHo, SFHx [53, 51], and LS220 with a nuclear incompressibility at saturation density of $K=220$ MeV [54]. Correspondingly, we follow Ref. [43] in our naming convention of the simulations, specifying the PNS mass and the EoS in the model names, e.g. 1.62-DD2. Our standard simulations include PNS convection and muons; those without convection are denoted by a suffix “-c” to their names, and those without muons by a suffix “-m”. In Table 1 we list all of the discussed simulations and the final post-bounce times $t_{\rm fin}$ when they were stopped.
@@
Figure 1 displays the time evolution of the luminosities $L_{\nu}(t)$ for all the neutrino (left panels) and antineutrino (right panels) flavors in units of $\mathrm{B/s}$ ($1~{}\operatorname{B}=10^{51}~{}\operatorname{erg}$) in the time interval [1,10] s for the 1.62 $M_{\odot}$ models, computed with the set of considered EoSs cases: DD2 in blue, SFHo in orange, SFHx in green, and LS220 in red. From this figure it is evident that the neutrino luminosities exhibit a change of their slope in the time interval considered. Since we expect a power-law behavior of the neutrino luminosities in the early cooling phase and we also want to enhance the visibility of the change in the steepness of the luminosity decline at later times, we take inspiration from Ref. [40] and show in Fig. 2 the product of the post-bounce time and the luminosities, $t\,L_{\nu}(t)$, for all of the neutrino (left panels) and antineutrino (right panels) flavors in the same time interval and for the same simulations as in Fig. 1. One witnesses an interesting difference of the results for the employed EoSs. For DD2, SFHo and SFHx EoSs and all of the neutrino species the quantity $t\,L_{\nu}$ increases as a power-law at $t\gtrsim 1~{}\operatorname{s}$ to peak around $t\approx 4~{}\operatorname{s}$, followed by a steep suppression at later times. This behavior produces a prominent knee at about 5–6 s after bounce. In contrast, for LS220 $t\,L_{\nu}$ reaches its peak earlier, namely at $t\lesssim 3~{}\operatorname{s}$, and it features a slower decline later on. Since for each model the luminosities of all heavy-lepton neutrino species show a similar evolution, we will focus exclusively on the $\bar{\nu}_{\mu}$ signal as representative of all non-electron neutrino species, unless otherwise noted.
@@
For each case shown in Figs. 2, 3, and 4, the quantity $t\,L_{\nu}$ has a peak at $t_{\nu_{i},\,\rm max}\lesssim 6$ s, when we define ${L_{\nu_{i},\,\rm max}\equiv L_{\nu_{i}}(t_{\nu_{i}\,\rm max})}$. The corresponding values of $t_{\nu_{i},\,\rm max}$ are listed in Table 2 for neutrinos and in Table 3 for antineutrinos in Appendix A. In Table 1 we provide the values of the quantity
@@
which is a measure for how long our simulations followed the late-time luminosity decline until they were stopped at $t_{\rm fin}$. For our benchmark PNS models (all models including muons and convection) $X_{\nu_{i}}^{\rm fin}\lesssim 0.15$ holds for all neutrinos and antineutrinos (the largest value is $X_{\nu_{e}}^{\rm fin}=0.123$ for 1.62-LS220). In order to adopt a well-defined, common final time for all of the neutrino species and simulations when fitting the luminosities in the following, we cut the data at an instant $t_{\nu_{i},\rm c}$ when $X_{\nu_{i}}^{\rm c}\equiv t_{\nu_{i},\rm c}L_{\nu_{i},\rm c}/(t_{\nu_{i},\rm max}L_{\nu_{i},\rm max})=0.15$. In addition to the values of $t_{\nu_{i},\rm max}$ we also provide those of $t_{\nu_{i},{\rm c}}$ for all neutrino species and simulations in Tables 2 and 3 in Appendix A. It can be seen that for a fixed EoS and neutrino species, the values of both of these quantities increase with the PNS mass. In contrast, given the PNS mass and EoS, $t_{\nu_{i},\,\rm max}$ and $t_{\nu_{i},{\rm c}}$ do not vary much among the different kinds of neutrinos and antineutrinos (e.g., for 1.62-DD2 $t_{\nu_{i},{\rm max}}\approx 4\,\operatorname{s}$ and $t_{\nu_{i},{\rm c}}\approx 8\,\operatorname{s}$ for all species). However, at fixed PNS mass, $t_{\nu_{i},\,\rm max}$ and $t_{\nu_{i},{\rm c}}$ feature a strong dependence on the considered EoS. In particular, $t_{\nu_{i},{\rm c}}\lesssim 10\,\operatorname{s}$ for all the benchmark models with DD2, whereas SFHo and SFHx lead to slightly larger values of $t_{\nu_{i},{\rm c}}$, with $t_{\nu_{i},{\rm c}}\lesssim 12.5\,\operatorname{s}$ for all models, and in the cases with LS220 we find the highest values of $t_{\nu_{i},{\rm c}}$, with $t_{\nu_{i},{\rm c}}\lesssim 15\,\operatorname{s}$ for the largest PNS mass.
@@ III Fit functions
where $C$, $\alpha$, $\tau$ and $n$ are free parameters and $t$ and $\tau$ are measured in seconds. [^4] We define the general parameters as $C$, $\alpha$, $\tau$ and $n$ and we use $C_{\nu_{i}}$, $\alpha_{\nu_{i}}$, $\tau_{\nu_{i}}$ and $n_{\nu_{i}}$ when referring specifically to a neutrino species $\nu_{i}$. In Eq. (2) the parameter $C$ is a normalization constant, $\alpha$ describes the power-law behavior at early cooling time deviating from the simple $t^{-1}$ behavior, $\tau$ is a characteristic cooling time for the exponential drop after the peak in $t\,L_{\nu}$, with $n$ representing the strength of the suppression at late times. We mention here that at the moment there is no straightforward argument explaining the power-law behavior at early cooling time, given that PNS convection plays an important role and it cannot be treated in simple ways. As an example, we show for the $1.62~{}M_{\odot}$ model and all EoSs the time evolution of $t\,L_{\nu_{i}}$ from simulations (blue lines) and its best fit (orange) between $t=1$ s and $t_{\nu_{i},{\rm c}}$, for $\nu_{e}$ in Fig. 5, for $\bar{\nu}_{e}$ in Fig. 6 and for $\bar{\nu}_{\mu}$ in Fig. 7. For DD2, SFHo and SFHx the fit well reproduces the results from the simulations in the considered time interval, with best-fit parameters in a similar range. For each EoS, the highest values of the fit parameters $C$ and $\tau$ are obtained for $\bar{\nu}_{\mu}$, while $\nu_{e}$ show the lowest ones. A similar behavior is found for the other PNS masses.
@@
The fit for models with EoS LS220 shows worse agreement with the data from simulations (see the lower-right panels in Figs. 5, 6, and 7). In particular, for all neutrino species the fit overestimates the luminosity at $t\approx 1\,$s and exhibits a lower value of the peak, whereas it well reproduces the luminosity at later times. In this case, the values of the best-fit parameters are in a completely different range compared to those of the models with the other EoSs. In particular, for each neutrino species $C\approx\mathcal{O}(100)~{}\operatorname{B}/\operatorname{s}$ and $\alpha<0$, whereas $\tau$ and $n$ adopt much smaller values than for the other considered EoSs, with $\tau\lesssim 0.1\,$s and $n\lesssim 0.5$, due to the more shallow decline of the luminosity at late times. The differences in the time dependence of the neutrino luminosities between models with different EoSs can be understood by the different behavior of the nuclear symmetry energy as a function of baryon density for the considered EoS. In the LS220 EoS, the symmetry energy exhibits a steeper increase with baryon density than in all other cases. As discussed in Ref. [42], a larger positive derivative of the symmetry energy with baryon density can lead to suppressed convection in the PNS mantle at high densities and low electron fraction. This effect happens in the simulations with the LS220 EoS after about 3 s and gradually quenches PNS mantle convection, thus delaying subsequent PNS neutrino cooling due to reduced neutrino luminosities. In contrast, PNS convection continues to be active in a spatially more extended region including the PNS mantle in simulations with the other three EoSs, for which reason convectively enhanced neutrino luminosities are maintained for a longer period of time until a late, steep decline follows when the PNS has deleptonized and cools off. These differences explain the early kink in the neutrino luminosities for models with the LS220 EoS, whereas a prominent knee-like shape of $L_{\nu}(t)$ can be witnessed for the DD2, SFHo, and SFHx models (see Fig. 1). These differences motivate us to define DD2, SFHo, and SFHx as members of an EoS-class that we call “Class A”, whereas LS220 is a representative of a “Class B”. The parameters of the symmetry energies for all the EoS cases used in our study are listed in Table 12 in Appendix F.